The sum of the digits of a two-digit number is 14. If 36 is subtracted from the number, the resulting number is the original number with its digits interchanged. Find the original number.
Answer by mananth(16949)

let the numbers be x & y x in the tens place.
x+y =14................(1)
--
The number is 10x+y
subtract 36
10x+y-36
new number = 10y+x
10x+y-36=10y+x
9x-9y=36
/9
x-y=4.................(2)
add (1) & (2)
2x=18
/2
x=9
if x= 9 y= 5
the number is 95